What is the least natural number which is larger than 100 and which leaves the remainder, R = 12 when it is divided by 19.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

100 = 19 × 5 + 5

As 5 + 7 = 12
the required number is 100 + 7 = 107

Thanks